OpenStudy (anonymous):

Solve the system of equations. Show work please, to help me understand y = 2x + 5 -2x + y + 5 = 0

OpenStudy (anonymous):

-2x+2x+5+5=0 -4x+10=0 after that you can no longer solve fo x

OpenStudy (sepeario):

y = 2x + 5 y = 0 - 5 + 2x 2x + 5 = 0 - 5 + 2x Can you work it from there? I'm still kind of learning this, I'm not really sure. You might need someone better with this to help...

jimthompson5910 (jim_thompson5910):

-2x + y + 5 = 0 -2x + 2x + 5 + 5 = 0 ... replace y with 2x+5 (since y = 2x+5) 0x + 10 = 0 0 + 10 = 0 10 = 0 this last equation is false, so there are no solutions